OpenStudy (ronaldo7):

2x² -7x += 15 Help Please

OpenStudy (ronaldo7):

Do you use the quadratic formula

OpenStudy (lochana):

2x^2 - 7x += 15 means 2x^2 - 7x = 2x^2 - 7x + 15

OpenStudy (lochana):

is it += or just =

OpenStudy (lochana):

I see. it is =.

OpenStudy (lochana):

no I don't use any formula

OpenStudy (ronaldo7):

Im confused what

OpenStudy (lochana):

in your original post, you have typed 2x² -7x "+=" 15 .

OpenStudy (ronaldo7):

Oh woops

OpenStudy (lochana):

that is why I said that

OpenStudy (ronaldo7):

its suppose to be 2x² -7x = 15

OpenStudy (ronaldo7):

sorry

OpenStudy (lochana):

2x² -7x -15= 0 2x^2 + (10x - 3x) -15=0 (2x^2 + 10x) - (3x +15)=0 2x(x + 5) -3 (x +5)=0 (2x-3)(x+5)=0

OpenStudy (ronaldo7):

So you use factoring

OpenStudy (lochana):

sorry. I made a mistake there

OpenStudy (lochana):

2x² -7x -15= 0 2x^2 + (-10x + 3x) -15=0 (2x^2 - 10x) + (-3x +15)=0 2x(x - 5) -3 (x -5)=0 (2x+3)(x-5)=0

OpenStudy (ronaldo7):

hold on im solving it all over again

OpenStudy (lochana):

yeah. I messed it:)

OpenStudy (ronaldo7):

I got (x-5)(2x +3) so the same thing

OpenStudy (lochana):

you need to make -7x into 2 parts so that you can pull out common factors like I did before

OpenStudy (ronaldo7):

then you set them both to zero and solve right?

OpenStudy (lochana):

yes. that's right.

OpenStudy (lochana):

and do you know why?

OpenStudy (ronaldo7):

Would you have to check the answers since its a radical equation

OpenStudy (ronaldo7):

Yeah I know why because to solve for y you must plug in 0 for x

OpenStudy (ronaldo7):

I mean to solve for x you plug in 0 for y

OpenStudy (lochana):

because. in order to be 0 of a multiplication of 2 terms happens only when one of them is equal to zero. so we argue that if (x-5)(2x +3) =0 (x-5) or (2x+3) has to be zero

OpenStudy (ronaldo7):

Oh yeah and that but would you ahve to check your answer since its a radical equation

OpenStudy (lochana):

are you still here?

OpenStudy (lochana):

2x² -7x -15= 0 2x^2 + (-10x + 3x) -15=0 (2x^2 - 10x) + (3x -15)=0 2x(x - 5) +3 (x -5)=0 (2x+3)(x-5)=0

OpenStudy (lochana):

today is a bad day:D